For the chlorinated hydrocarbon dichlorobutane how many electrons are being used to form covalent bonds? How many electrons are being shared(Do not include core electrons)
1
Expert's answer
2020-11-09T14:10:21-0500

C4H8Cl2

Number of C-C bonds = 3

Number of C-H bonds = 8

Number of C-Cl bonds = 2

Total bonds = 13

One bond contains 2 e-

13 bonds contain "13 \times 2 e^- = 26 \;e^-"

26 electrons are being used to form covalent bonds.

The same number of electrons (26) are being shared.